### v1

# Probing the cosmic sterile-neutrino background with IceCube

Abstract
In this paper, we take a close look at the interaction between the TeV–PeV energy astrophysical neutrinos and a hypothetical cosmic sterile-neutrino background. These interactions yield absorption features, also called “dips”, in the astrophysical neutrino spectrum, which are studied using the deposited energy distribution of high-energy starting events (HESE) in the IceCube detector. We improve upon the previous analysis by including the effects of regeneration and a realistic source distribution on the propagation of astrophysical neutrinos. We use the latest 7.5-year HESE dataset and include the observation of Glashow resonance in our analysis. We evaluate the impact of these dips on the inferred spectral index and overall normalization of the astrophysical neutrinos. We find a mild preference for dips in the 300–800 TeV range, and the best-fit parameters for the mass of sterile-neutrino and the mediator are 0.5 eV and 23 MeV, respectively. We find that the inclusion of these absorption features lowers the spectral index of astrophysical neutrinos to $2.60^{+0.19}_{-0.16}$. The lower spectral index can reduce the disagreement with the Northern Tracks sample but requires dedicated analysis. We also forecast the event spectrum for IceCube-Gen2 for the two different fits.
I Introduction
The IceCube neutrino observatory has detected neutrinos of astrophysical origin in the TeV–PeV range in several detection channels. The latest 7.5 years (2635 days) high energy starting events (HESE) sample consists of 102 events, of which 60 events have reconstructed deposited energy above 60 TeV Abbasi et al. (2021). The IceCube experiment has also observed a Glashow Resonance (GR) event Glashow (1960) in the 4.6 years PeV energy partially contained events (PEPE) sample Aartsen et al. (2021a). IceCube also maintains a catalog of through-going muon-neutrinos from the northern hemisphere (henceforth called Northern Tracks). A large fraction of the Northern Tracks are below 100 TeV and originate via atmospheric neutrinos. The highest energy events from this sample have astrophysical origin.
It is abundantly clear that these neutrinos have an extra-galactic origin and are likely to originate in sources capable of accelerating cosmic rays to EeV-scale energies. The neutrinos are produced in subsequent interactions of these ultra-high energy cosmic rays with ambient gas or radiation. At the time of writing, only a handful of such sources have been confirmed Aartsen et al. (2018a, b, 2020); Abbasi et al. (2022a). If the neutrinos are produced via interactions with ambient gas, i.e. in proton-proton ($pp$) interactions, the neutrinos are expected to have a power-law spectrum Margolis et al. (1978); Stecker (1979); Kelner et al. (2006). In the baseline scenario, the neutrinos oscillate and reach flavor equilibrium before they arrive to the Earth Lipari et al. (2007); Choubey and Rodejohann (2009); Mena et al. (2014, 2014); Aartsen et al. (2015). One usually assumes that an isotropic flux of astrophysical neutrinos is incident on the Earth with a total differential all-flavor neutrino-plus-anti-neutrino spectrum given by
$\dfrac{d\Phi_{6\nu}}{dE_{\nu}}=\Phi_{\texttt{astro}}\,\left(\frac{E_{\nu}}{E_{0}}\right)^{-\gamma_{\texttt{astro}}}\,\cdot 10^{-18}\,\text{GeV}^{-1}\,\text{cm}^{-2}\,\text{s}^{-1}\,\text{sr}^{-1},$ (1)
which is characterized by two free parameters: a spectral index ($\gamma_{\texttt{astro}}$) and the normalization scale ($\Phi_{\texttt{astro}}$). The reference energy scale $E_{0}$ is chosen to be 100 TeV. If the neutrinos are produced via photo-hadronic interactions, i.e. in proton-gamma ($p\gamma$) sources, the energy spectrum is expected to be “bump-like” Kelner et al. (2006); Mucke et al. (2000); Hummer et al. (2010); Fiorillo and Bustamante (2023). Realistically, both $pp$ and $p\gamma$ sources are expected to contribute. In this work, we will assume that the $pp$ sources dominate and the diffuse flux of astrophysical neutrinos is described by a single power-law (SPL), as in Equation 1.
Are the events observed by IceCube consistent with the assumption of SPL? To a large extent, yes. However, there are some tantalizing discrepancies that have motivated a large body of work in the last decade. Some of the explanations are based on dark matter Feldstein et al. (2013); Esmaili and Serpico (2013); Bai et al. (2016); Fong et al. (2015); Rott et al. (2015); Esmaili et al. (2014); Murase et al. (2015); Anchordoqui et al. (2015); Bhattacharya et al. (2017a, b); Hiroshima et al. (2018); Chakravarty et al. (2020), resonances beyond standard model Anchordoqui et al. (2006); Barger and Keung (2013); Dutta et al. (2015); Dey and Mohanty (2016); Dev et al. (2016); Mileo et al. (2016); Chauhan et al. (2018); Dey et al. (2018); Bečirević et al. (2018); Babu et al. (2020); Dey et al. (2021), or non-standard/secret neutrino interactions Ng and Beacom (2014); Ioka and Murase (2014); Ibe and Kaneta (2014); Blum et al. (2014); Araki et al. (2015); Cherry et al. (2014); Kamada and Yu (2015); DiFranzo and Hooper (2015); Araki et al. (2016); Rasmussen et al. (2017); Jeong et al. (2018); Mohanty et al. (2019); Pandey et al. (2019a, b); Bustamante et al. (2020); Creque-Sarbinowski et al. (2021); Mazumdar et al. (2022); Fiorillo et al. (2020a, b); Carpio et al. (2023); Esteban et al. (2021). One such discrepancy is the inferred value of the spectral index in two independent datasets,
[…]
We begin with the ansatz that including “dips” in the neutrino flux will result in a lower spectral index from the HESE sample and thus relax the tension. Such dips are a consequence of interaction between astrophysical neutrinos and cosmic background. If the interaction is through a s-channel process, then the absorption feature is narrowly located around the resonance energy $E_{\nu}^{\rm res}=M_{X}^{2}/2m_{b}$, where $M_{X}$ is the mass of the mediator and $m_{b}$ is the mass of the background particle. If the interaction is t-channel, then a broad spectrum feature is expected Mohanty et al. (2019). In Ref. Esteban et al. (2021), the effect of non-resonant contribution to the process was found to be significant for larger couplings, and we have included them in our analysis. With only standard model interactions, the dip will occur for neutrinos with energy EeV Weiler (1982); Brdar et al. (2022). However, if there are non-standard/secret interactions mediated by a MeV-scale boson, then the dips can occur in the TeV–PeV range Ng and Beacom (2014). In this paper, we study the scenario proposed in Ref. Chauhan and Mohanty (2018) where the dip is caused by interactions between astrophysical neutrinos and a cosmic sterile-neutrino background.
Figure 1: The test statistic (see Equation 5) for the single power-law fit of 7.5 years HESE sample Abbasi et al. (2021) is shown in blue. In orange, we show the test statistic after including absorption features and Glashow resonance. The horizontal grey dashed lines are used to denote the 68% C.L. and 90% C.L. intervals. We have shown the results from the 11-year Northern Tracks sample Stettner (2020) in green.
The light sterile neutrino is a well-motivated extension of the standard model and has been the subject of investigation in a number of experiments. Currently, there is evidence in support of  Anselmann et al. (1995); Hampel et al. (1998); Aguilar-Arevalo et al. (2009, 2010, 2001); Mention et al. (2011); Aguilar-Arevalo et al. (2018); Serebrov et al. (2021) as well as against Abratenko et al. (2023) the sterile neutrino. For a recent review, we refer to Ref. Dasgupta and Kopp (2021). The presence of a fully thermalized sterile neutrino is in conflict with the measurement of $N_{\text{eff}}$ from cosmological probes Hamann et al. (2011); Cooke et al. (2018); Aghanim et al. (2020). If the sterile neutrino has self-interactions via a MeV scale vector mediator, their recoupling can be delayed, and the $N_{\text{eff}}$ constraints can be avoided Hannestad et al. (2014); Dasgupta and Kopp (2014). The case for pseudo-scalar mediator is studied in Ref. Archidiacono et al. (2015). Despite the early success, the self-interacting sterile neutrino scenario has been tested against many cosmological observations Saviano et al. (2014); Mirizzi et al. (2015); Forastieri et al. (2017); Choi et al. (2018); Song et al. (2018). It was shown in Ref. Chu et al. (2018) that the basic scenario is in conflict with bounds on the sum of neutrino mass, and some resolutions have been provided as well. One must note that these results depend on the choice of background cosmology as $\Lambda$CDM, and the bounds will change if other scenarios are considered Roy Choudhury and Naskar (2019); Roy Choudhury and Choubey (2018); Roy Choudhury and Hannestad (2020). Recent analyses of large-scale structure data have also indicated the preference for self-interacting neutrinos He et al. (2024); Camarena et al. (2023). With this as motivation, we ask if the dip due to a cosmic sterile-neutrino background can be probed in IceCube and whether the inclusion of these features leads to an appreciable decrease in the inferred spectral index. The main result of this paper is shown in Figure 1. We find that the systematic inclusion of the dips results in,
[…]
which is in better agreement with the Northern Tracks sample. As discussed below, the errors are underestimated due to the limitations of our analysis. For a quantitative assessment of the relaxation in the tension between the two datasets, one must evaluate the impact of dips in the spectrum on the Northern Tracks sample. However, this exercise is beyond the scope of our paper. We have shown that an ‘SPL with dips’ is a viable fit to the present HESE sample and results in a lower spectral index than an SPL-only explanation. One must note that our conclusion also extends to other origins of the dip(s).
The paper is organized as follows. In Section II, we discuss the model, propagation of astrophysical neutrinos, our estimate of the HESE event rate, and the likelihood. In Section III, we provide the main results and end with a short summary in Section IV.
II Model description and methods
In this work, we follow the model outlined in Ref. Dasgupta and Kopp (2014), which is briefly described below. The sterile neutrinos are assumed to have self-interactions mediated by a new MeV-scale vector boson. The interaction Lagrangian is given by,
$-\mathcal{L}_{\rm int}=g_{X}\bar{\nu}_{s}\gamma_{\mu}\frac{1}{2}(1-\gamma_{5})\nu_{s}X^{\mu}\,,$ (2)
where $X^{\mu}$ is the mediator and $g_{X}$ is the self-interaction coupling constant. In this work, we fix the coupling strength as $g_{X}=0.1$, which is sufficiently large for the re-coupling of sterile neutrinos Chu et al. (2015). The mass-eigenstates and the flavor-eigenstates are related by a $4\times 4$ rotation matrix characterized by six mixing angles. The three mixing angles from the active sector are fixed to their measured values from Ref. Esteban et al. (2020), and the remaining three angles are free parameters of the model. Due to this mixing, all four mass-eigenstates interact via the new mediator, and interactions between propagating mass-eigenstates and background neutrinos are possible Chauhan and Mohanty (2018). In this simplified model, we have five new physics parameters – three mixing angles ($\theta_{14}\,,\theta_{24}\,,\theta_{34}$), mass of the sterile neutrino ($m_{s}$), and mass of the mediator ($M_{X}$). In a UV-complete model, additional degrees of freedom may be required for generating the mixing and mass of the gauge boson and for anomaly cancellation. The alternative model with a pseudoscalar mediator is cosmologically more viable Dasgupta and Kopp (2021), but has additional constraints from meson decays Fiorillo et al. (2020a). We anticipate that similar bounds will exist on vector mediators, but their evaluation is beyond our scope. Our main conclusion, i.e., the spectral index is lowered because of the dips, does not depend on the details of the model causing the dips in the spectrum.
The introduction of this new interaction results in scattering between astrophysical neutrinos and the cosmic background of active and sterile neutrinos. Consequently, we expect dips and bumps in the flux of astrophysical neutrinos, which is otherwise assumed to have a single-power law behavior. The Boltzmann equation for the time evolution of the comoving number density of the astrophysical neutrinos per unit energy ($\tilde{n}_{i}=\partial\Phi_{i}/\partial E_{\nu}$) is given by Ng and Beacom (2014)
$\frac{\partial\tilde{n}_{i}(t,E_{\nu})}{\partial t}$ $=\frac{\partial}{\partial{E_{\nu}}}\left[H(t)\,E_{\nu}\,\tilde{n}_{i}(t,E_{\nu})\right]+\mathcal{L}_{i}(t,E_{\nu})-\tilde{n}_{i}(t,E_{\nu})\sum_{j}n^{t}_{j}\sigma_{ij}(E_{\nu})$
[…]
In this equation, the first term on the right-hand side accounts for the cosmological expansion. The second term, called the source term, accounts for the production of high-energy neutrinos due to astrophysical sources. In this work, we assume that the TeV–PeV neutrinos originate in $pp$-sources and following Ref. Esteban et al. (2021), the source term is taken to be a single power law with normalization proportional to the star-formation rate. The third term accounts for the interaction with background neutrinos, and the last term accounts for the regeneration of lower-energy neutrinos after interactions. These equations have been solved using the publicly available code nuSIprop Esteban et al. (2021), which solves the Boltzmann equations for standard three-flavor neutrinos in the presence of non-standard interaction. We have extended the code to include the sterile neutrino. These are improvements on the work done in Ref. Chauhan and Mohanty (2018) where the source term was ad-hoc, and regeneration was ignored.
To estimate the rate of high-energy starting event (HESE) in IceCube, we use the Monte Carlo sample of simulated events [^1] https://github.com/icecube/HESE-7-year-data-release/ from Ref. Abbasi et al. (2021). The input parameters used to calculate the HESE event spectrum are tabulated in Table 1, and a description of these parameters can be found in Ref. Abbasi et al. (2021). Of these, we only consider the astrophysical neutrino spectral index ($\gamma_{\texttt{astro}}$) and the normalization scale ($\Phi_{\texttt{astro}}$) as variables and fix other parameters to their typical values. We assume that the new physics does not affect the fluxes of conventional and prompt atmospheric neutrinos, as well as atmospheric muons. A detailed calculation where all parameters are optimized is computationally very expensive and beyond the scope of this work. The event spectrum and the subsequent log-likelihood are obtained by reweighting the Monte Carlo events from astrophysical neutrinos. As we also scan over the mixing angles, the reweighting accounts for the flavor dependence of the interactions.
Table 1: The set of physical and nuisance parameters that determine the HESE event rate are tabulated. We have shown the best-fit parameters for single power law (SPL) obtained in Ref. Abbasi et al. (2021) and the values/range of parameters used in this work. For a description of these variables, see TABLE VI.1 in Ref. Abbasi et al. (2021)
Parameter SPL best-fit This Work
$\gamma_{\texttt{astro}}$ 2.874 (2,4)
$\Phi_{\texttt{astro}}$ 6.365 (1,10)
$\Phi_{\texttt{conv}}$ 1.006 1.000
$\Phi_{\texttt{prompt}}$ 0.0 0.0
$\Phi_{\texttt{muon}}$ 1.187 1.00
$\Delta\gamma$0.053 -0.050
$\nu/\overline{\nu}$ 0.998 1.000
$a_{s}$ 1.001 1.000
$\varepsilon_{\rm dom}$ 0.952 0.990
$\varepsilon_{\rm h.o.}$0.055 0.000
$K/\pi$ 1.000 1.000
We follow the analysis of Ref. Abbasi et al. (2021) closely and present our results in the frequentist statistical methodology. We report the new physics parameters that are most likely to explain the HESE data and compare it with the single power-law. The best-fit parameters and the errors are obtained using the profile likelihood,
~$\tilde{\mathcal{L}}^{\texttt{profile}}(\bm{\theta})=\max_{\bm{\eta}}\mathcal{L}(\bm{\theta,\eta})\cdot\Pi(\bm{\theta,\eta})$ (4)
where $\bm{\theta}$ is the set of parameters of interest and $\bm{\eta}$ is the set of nuisance parameters. The likelihood function used is the effective likelihood defined in Ref. Argüelles et al. (2019). The constraint $\Pi(\bm{\theta,\eta})$ is constant for parameters for which no external data is available. We have used the Python implementation of SLSQP algorithm to minimize the negative of the log of the function and thus obtain the best-fit point $\hat{\bm{\theta}}$. The model parameter test statistic is defined as,
$\text{TS}(\bm{\theta})=-2\,\log\,\left(\dfrac{\tilde{\mathcal{L}}^{\texttt{profile}}(\bm{\theta})}{\tilde{\mathcal{L}}^{\texttt{profile}}(\hat{\bm{\theta}})}\right).$ (5)
The results are presented using Wilks’ theorem with appropriate degrees of freedom Wilks (1938). To validate our method, we first evaluate the 68 % C.L. and 90 % C.L. intervals for single-power law and compare them with the contours in Ref. Abbasi et al. (2021). We only scan over the astrophysical flux parameters and keep other parameters fixed to their best-fit values. We find that on expected lines, the best-fit point agrees very well, and the errors are underestimated.
The HESE sample only includes events for which the interaction vertex is inside the restricted fiducial volume. To enhance the sensitivity to Glashow resonance (GR) events, a larger fiducial volume is required, which is achieved by looking at PeV energy partially contained events (PEPE). As the Monte Carlo event sample contains the GR, we use a simple rescaling for exposure to estimate the GR event rate. We find that our estimates agree with the Extended Data Table 1 in Ref. Aartsen et al. (2021a). The likelihood function, i.e., $\mathcal{L}(\bm{\theta,\eta})$ in Equation 4, is modified to include GR event for which we assume a simple Poisson likelihood.
III Searching for dips in spectrum
In this section, we look at the range of new physics parameters that lead to significant dips in the HESE sample and evaluate their impact on the spectral index ($\gamma_{\texttt{astro}}$) and the normalization scale ($\Phi_{\texttt{astro}}$) of the diffuse astrophysical flux. In our analysis, we fix the values of the parameters related to the detector efficiency, cosmic ray flux, and atmospheric neutrinos (see Table 1). As a result of this approximation, the errors in this paper are underestimated. A detailed scan of all parameters is computationally intensive and beyond the scope of this work. The parameters of interest are,
$\bm{\theta}=\{\gamma_{\texttt{astro}},\,\Phi_{\texttt{astro}},\,M_{X},\,m_{s},\,\theta_{14},\,\theta_{24},\,\theta_{34}\}.$ (6)
The best-fit values of these parameters from our analysis is tabulated in Table 2. We fix the mixing angles to their best-fit values for the rest of the analysis. This reduces the computation time significantly and allows us to scan over relevant parameter space only. This approximation underestimates the 90% C.L. region, and the 68 % C.L. region is mostly unaffected.
Table 2: The best-fit values of the parameters of interest for single power law (SPL) obtained in Ref. Abbasi et al. (2021) and the result from our analysis is tabulated.
Parameter SPL best-fit This Work
$\gamma_{\texttt{astro}}$ 2.874 2.613
$\Phi_{\texttt{astro}}$ 6.365 5.461
$M_{X}$ 22.96 MeV
$m_{s}$ 0.50 eV
$\theta_{14}$ 0.52
$\theta_{24}$ 0.53
$\theta_{34}$ 0.41
Figure 2: Scan of the profile likelihood (see Equation 5) for the two new physics parameters: mass of the sterile neutrino ($m_{s}$) and mass of the vector boson mediator ($M_{X}$), is shown in orange. The darker (lighter) region denotes the 68% C.L. (90 % C.L.) interval. Note that for each point, the mixing angles are fixed to their best-fit values, and all other parameters are optimized. We also show two approximate curves that result in dips at 300 TeV and 800 TeV due to interaction with a cosmic sterile neutrino background. The region labeled as NSI (pink) indicated the parameter space for which dips occur due to interaction with the active neutrino background only and the dip due to sterile neutrino is outside the HESE analysis bound.
We first look at the space of the new physics parameters – sterile neutrino mass ($m_{s}$) and the vector mediator mass ($M_{X}$), and present our results in Figure 2. As expected [^2] The Figure 2 should be compared with Fig. 2 in Chauhan and Mohanty (2018), we find a large degenerate parameter space that results in dips in the event spectrum. We also find an island for $M_{X}=5-10$ MeV, which is largely independent of the sterile neutrino mass. In this region, the dip due to sterile neutrino is below the 60 TeV threshold of the HESE sample, and there is a dip between 300-800 TeV due to the heavier active neutrino. This is similar to the case of neutrino secret interaction ($\nu$SI), which is considered in Ref. Ng and Beacom (2014). We also highlight the $m_{s}=2.6$ eV that is hinted in recent results Serebrov et al. (2021); Abbasi et al. (2022b).
Figure 3: Scan of the profile likelihood for the two new physics parameters: astrophysical normalization and spectral index. The darker (lighter) regions show the 68 % C.L. (90 % C.L.) intervals. The region in blue is obtained using single power law, and the region in orange is obtained by including Glashow Resonance and dips due to sterile neutrino background. Note that other physical parameters are optimized for each scan point. As we do not optimize over all of the parameters (see text for details), the intervals are slightly smaller than the ones reported in Ref. Abbasi et al. (2021).
Figure 4: Left: We show the best-fit SPL flux from Ref. Abbasi et al. (2021) (blue) and the best-fit flux from our analysis (orange). The binned reconstructed flux from the HESE sample Abbasi et al. (2021) and PEPE sample is shown in black and red, respectively. Right: The deposited energy spectrum of HESE events (black) is shown with our event prediction using the best-fit parameters for SPL with dips (orange) and SPL-only (blue). The background from atmospheric neutrinos and muons is the same for both curves. The shaded grey region represents the 60 TeV threshold below which the backgrounds dominate the event rate.
We now look at the space of astrophysical neutrino flux parameters – spectral index ($\gamma_{\texttt{astro}}$) and the normalization scale ($\Phi_{\texttt{astro}}$). The scan of the profile likelihood is presented in Figure 3. We find that the inclusion of the dips leads to lower values of the spectral index. The inclusion of the Glashow resonance event also disfavors large values of the spectral index. This is one of the key results of the paper. At present, with only 7.5 years of data, there is a large overlap between the preferred regions for SPL and SPL with dips. However, the trend is clearly noticeable. In the future, with additional data from other neutrino telescopes, we anticipate that the overlap will be resolved. This was recently demonstrated in Ref. Fiorillo and Bustamante (2023), where the authors search for a bump-like feature in the spectrum. We also look at single parameter profile likelihood, and the results have been presented in Figure 1. It is clear that the inclusion of absorption features leads to a lowering of the spectral index and thus reduces the tension in the HESE and Northern Tracks samples. A future systematic study would also include the impact of absorption on the Northern Tracks sample.
In order to highlight the difference between the SPL best-fit and the best-fit parameters for SPL with dips, we show the neutrino flux prediction overlaid with the reconstructed flux from the IceCube collaboration in Figure 4. This plot only indicates of the position of dips in the neutrino spectrum. In our analysis, the likelihood is computed using the data from the deposited energy event spectrum. We have shown the predicted event spectrum from the two sets of best-fit parameters in Figure 4. As mentioned earlier, we assume that the atmospheric neutrinos are unaffected by the new physics, and the background predictions for both fits are the same. In a more detailed analysis, one may consider varying the atmospheric neutrino parameters. As there is a large overlap in the 68% C.L. interval of the two fits, we choose not to show the uncertainty bands in these figures.
Figure 5: Event rate prediction for 25 years of exposure for IceCube-Gen2. The orange curve represents events assuming the best-fit parameters of this work (see Table 2). The blue curve represents events for the SPL best-fit parameters. The blue-shaded region shows 68 % C.L. errors on the SPL event rate assuming Poisson distribution. The grey-shaded region represents the 200 TeV threshold for IceCube-Gen2 Aartsen et al. (2021b).
III.1 Forecast for IceCube-Gen2
IceCube-Gen2 is the planned upgrade of the IceCube neutrino observatory, which is expected to have $\sim$10 times larger exposure to TeV–PeV neutrinos Aartsen et al. (2021b). To estimate the event rate in IceCube-Gen2, we scale the weights according to the increased effective area. The effective area for cascades in IceCube-Gen2 is given in Ref. Aartsen et al. (2021b). However, the effective area for tracks in IceCube-Gen2 is not publicly available at the time of writing this paper. In this work, we will assume that the effective area is the same for cascades and tracks. We have evaluated the event rates using 25 years of IceCube-Gen2 exposure for the best-fit parameters from the SPL analysis and for the best-fit point from this work. The results are presented in Figure 5. The local significance of the dips can be tested with such large exposures. It must be noted that there are other proposed neutrino telescopes that will also be sensitive to the absorption features. A combined analysis will improve the discovery prospects Fiorillo and Bustamante (2023).
IV Summary and outlook
The IceCube neutrino observatory has played a crucial role in our understanding of astrophysical neutrinos in the last decade. It also provides an opportunity to study rare interactions over large distances. In this paper, we have taken a close look at the interaction of the astrophysical neutrinos with a hypothetical sterile neutrino background. The light sterile neutrino is well-motivated from short baseline experiments. However, there is a conflict with standard cosmology, and one of the resolutions is to allow self-interactions via a MeV-scale mediator. As a consequence, one expects the cosmic neutrino background to have a sterile neutrino component as well. The astrophysical neutrinos can interact with the sterile neutrino background, which leads to dips in the TeV–PeV neutrino energy spectrum. If the dips are significant, there is an observable impact on the HESE sample in IceCube.
We have improved on the previous analysis on multiple fronts. We have taken a realistic source distribution and included the effect of regeneration during the propagation. We also account for the recent observation of Glashow resonance in our analysis. To estimate the likelihood, we first compute the event rate and then compare with the observations. We find that the inclusion of these absorption features allows for a lower spectral index ($\gamma_{\texttt{astro}}=2.60^{+0.19}_{-0.16}$) when compared with the single-power law flux ($\gamma_{\texttt{astro}}=2.87^{+0.20}_{-0.19}$). Note that, due to the approximations in our analysis, our errors are underestimated. The lower spectral index is in better agreement with the Northern Tracks sample ($\gamma_{\texttt{astro}}=2.28^{+0.08}_{-0.09}$). However, a detailed study that accounts for the effect of these absorption features on the Northern Tracks sample is essential and left as future work. Currently, with only 7.5 years of HESE data, we find that there is a large overlap of parameter space, and one cannot statistically infer the existence of such dips. In the future, with a much larger neutrino telescope in the form of IceCube-Gen2 and perhaps other IceCube-scale experiments, the absorption features will be strongly tested.
Acknowledgments
The authors would like to thank Austin Schneider for helping with the HESE Monte Carlo simulation and Ivan Esteban for helping with nuSIprop. The authors would also like to thank Subhendra Mohanty and Ivan Esteban for their comments and suggestions on the manuscript. PP thanks Ranjan Laha for useful discussion. BC is supported by an OPERA grant and seed grant NFSG/PIL/2023/P3821 from BITS Pilani. PP acknowledges the support from the National Post-Doctoral Fellowship by the Science and Engineering Research Board (SERB), Department of Science and Technology (DST), Government of India (SERB-PDF/2023/002356) and the IOE-IISc fellowship program.